use rational exponents to write x^1/8*y^1/4*z^1/2 as a single radical expression.: use rational exponents to write x^1/8*y^1/4*z^1/2 as a single radical expression.
Answer by Mathtut(1309)

x^1/8*y^1/4*z^1/2 = (x*y^2*z^4)^1/8 = eighth root of (x*y^2*z^4)